# {textblock*}

### v2

@@ I Introduction
However, a newly released result on the primordial power spectrum from the Atacama Cosmology Telescope (ACT) DR6 [5, 6] reported a slightly larger spectral index. Combined with Planck data at large-to-intermediate scales (denoted as P-ACT) and together with CMB lensing and BAO data from DESI [7] (denoted as P-ACT-LB), the analysis yields a new constraint on the spectral index: $n_{s}=0.9743\pm 0.0034$ (68% CL) at the CMB pivot scale $k_{\mathrm{CMB}}=0.05~\mathrm{Mpc}^{-1}$ [5]. This result appears to disfavor the $R^{2}$ inflation as illustrated in Fig. 5. In addition, the error bar for the running of the spectral index $\alpha_{s}$ is tightened in the P-ACT-LB result, which suggests a slight positive running of the spectral index $\alpha_{s}=0.0062\pm 0.0052$ (68% CL) [6]. The constraint on $\alpha_{s}$ is not consistent with most of the single-field models which involve a convex potential, including the $R^{2}$ model. Several works [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] discussed extensions to the single-field $R^{2}$ model that may resolve this inconsistency.
@@
A crucial point in our model is the presence of a new boson field $\chi$ which has an $Z_{2}$ invariant potential where the $\chi$ is odd under the discrete symmetry. We identify it with the $Z_{4}$ symmetry whose breaking generates the Majorana masses $M_{i}$ for the right-handed neutrinos $N_{Ri}~(i=1-3)$  [34]. Here, the $\chi$ and $N_{Ri}$ have the $Z_{4}$ charges +2+2+ 2 and 1-1- 1, respectively. This $Z_{4}$ is free from the Dai-Freed anomaly [35] and hence it is a perfect symmetry [36, 37] in the standard model. The breaking scale of the $Z_{4}$ is $\langle\chi\rangle\simeq 10^{12}$ GeV which gives the Majorana masses $M_{i}\sim g_{i}\langle\chi\rangle$. Those masses are very much consistent with the observed small masses for the neutrinos [38, 39, 40, 41].
II The $\chi$-extended $R^{2}$ model
We consider the action in which $R^{2}$ gravity is non-minimally coupled to a scalar field $\chi$ [20, 10, 21, 22],
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}f(R,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)\bigg{]},$ (1)
@@ II The $\chi$-extended $R^{2}$ model
where $v=\pm m/\sqrt{\lambda}$ is the vacuum expectation value of $\chi$, and $m$ is the mass of $\chi$. Notice that the above action is invariant under the $Z_{4}$ symmetry under which the $\chi$ transforms to $-\chi$. The presence of $|\chi|$ is not a serious problem, since we consider it to be the low-energy action in an effective field theory (EFT). In fact, the $|\chi|$ dependent terms often appear at the EFT level when such a boson field couples to non-abelian gauge theories. (See [43] for example.)
It is well-known that the $R^{2}$ term introduces a scalar degree of freedom to the theory. Introducing an auxiliary field $\psi$, the action (1) is equivalent to,
$S_{\psi}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}RF(\psi,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi$ (2)
$-V(\chi)-U(\psi)\bigg{]},$
@@
$F(\psi,\chi)=1+\frac{\psi}{6M^{2}}-\frac{\xi}{M_{\mathrm{pl}}^{2}}(|\chi|-\chi_{0})^{2}$ (3)
$U(\psi)=\frac{M_{\mathrm{pl}}^{2}\psi^{2}}{12M^{2}}$ (4)
To bring the above action into the Einstein frame, we perform the conformal transformation $\tilde{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ [44] for the choice of $\Omega^{2}=F$, where the $\Omega^{2}$ is the conformal factor, and a tilde represents quantities in the Einstein frame. Under transformation, the action (1) can be written in the Einstein frame,
$S_{\mathrm{E}}=\int d^{4}x\sqrt{-\tilde{g}}\left[\frac{M_{\mathrm{pl}}^{2}}{2}\tilde{R}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right.$ (5)
$\left.-\frac{1}{2}F^{-1}\tilde{g}^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-U(\phi,\chi)\right],$
@@
    Given the large initial value of $\phi$, the spectator field $\chi$ acquires an effective mass approximately given by $m_{\chi}^{2}\approx 3M^{2}\xi(1+\mathcal{O}(F(\phi)^{-1}))$. We assume that the initial value of $\chi$ is close to its potential minimum ±$\pm\chi_{0}$ at the beginning of inflation. During this stage, when $m_{\chi}^{2}>9H^{2}/4\approx 9M^{2}/16$, or equivalently $\xi>3/16$, $\chi$ is considered a heavy field and behaves as an underdamped oscillator. In contrast, when $m_{\chi}^{2}<9H^{2}/4$, $\chi$ is light and behaves as an overdamped oscillator; note that in this case, $\phi$ remains the lightest field. In each of the cases, the scalar field $\chi$ is settled down to one of its potential minimum ($\chi_{0}$ or $-\chi_{0}$, $\chi_{0}$ is taken for the computations in the following session.) before this stage of inflation reaches an end.
@@
    $U(\phi,\chi)\approx\frac{m^{4}}{4\lambda}-\frac{1}{2}m^{2}\chi^{2}$ (7)
    As shown above, the boson field $\chi$ becomes unstable $(m_{\chi}^{2}\approx-m^{2}<0)$. Therefore, the inflaton dynamics begin to turn toward the ±$\pm\chi$ direction. It is worth mentioning that $\chi$ field starting from a initial condition away from 0 steps down the potential during the first stage of the $R^{2}$ inflation and hence the $Z_{4}$ symmetry is already broken for the background trajectory, or more precisely,inflation selects a vacuum branch during the early stage (see in Fig. 1). When fixing $|\chi_{0}|^{2}\gg\langle\delta\chi^{2}\rangle$  [45, 46], the overproduction of domain walls (DWs) can be avoided, which usually happens in hybrid inflation models with a discrete $Z_{2}$ symmetry.
  • St-2: Effective single field $\chi$ inflation. The second stage of inflation starts along $\chi$ direction from ±$\pm\chi_{0}$ (the sign will depend on the initial condition for $\chi$), with the heavy field $\phi$ stabilized in its potential minimum. We may observe that $\phi$ is still evolving during this stage, but it is simply due to the varying $\chi$ causing the movement of the potential minimum. The time evolution of $\phi$ is not violating the effective single field slow condition, since the isocurvature is heavy and decays out very rapidly. The second stage of inflation ends as an effective single field inflation of $\chi$. To have a sizable number of efolds during St-2, $v\gg\chi_{0}$ is required. Combining with the non-DW condition, the constraints for the parameters is given
    $m^{2}/\lambda\gg\chi_{0}^{2}\gg M^{2}/(64\pi^{2}).$ (8)
@@
Thus, by varying the coupling constant $\xi$ from 0 to values 3/16$\geq$ 3/16≥ 3 / 16, the model allows the contribution from $\chi$ to the adiabatic power spectrum to evolve from nearly scale-invariant ($k^{0}$) to steeply blue-tilted ($k^{3}$), as explained detailedly in the following section (also shown in [22]).
@@
III Semi-analytical power spectrum
@@
For simplicity, we assume the transition between the two inflationary phases occurs instantaneously.
@@
$\mathcal{R}=\delta N=\delta N_{1}+\delta N_{2},$ (9)
[…]
@@
If we conduct the $\delta N$ calculation nonlinearly, one may observe a large deviation of the probability distribution function of $\mathcal{R}$ from Gaussianity. Such non-Gaussian features are potentially detectable by upcoming observations. We leave this for future work.
@@
$\delta N=\left(\frac{\partial N_{1}}{\partial\phi}\delta\phi_{k}\right)_{n_{k}}+\left(\frac{\partial N_{2}}{\partial\chi}\delta\chi_{k}\right)_{n_{\star}},$ (10)
@@
$\delta\chi_{k}(n_{\star})\approx\delta\chi_{k}(n_{k})e^{-a(n_{\star}-n_{k})/2},$ (11)
@@
$a$ $\equiv\mathrm{Re}\left(3-3\sqrt{1-\frac{16m^{2}_{\chi}}{9M^{2}}}\right)$ (12)
@@
during the St-1, which equals the maximum 333 for the heavy $\chi$ case and smaller when decreasing $\xi$.
Figure 3: The $n_{s}-\alpha_{s}$ relation shown in Eq. (14) for different $M_{\mathrm{PBH}}$s together with the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). To plot the figure, we use $\mathcal{P}_{\mathcal{R}}^{\mathrm{max}}/\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=10^{7}$ in order to have $f_{\mathrm{PBH}}=1$. The hollow black dots refer to cases that $a=3$, or equivalently $\xi\geq 3/16$ when the small-scale amplification is well isolated from the large scales.
@@
$\mathcal{P}_{\mathcal{R}}(k)=\mathcal{P}_{\mathcal{R}}^{R^{2}}(k)+\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}\left(\frac{k}{k_{1}}\right)^{a}.$ (13)
@@
$n_{s}\approx 1-\frac{2}{N_{1}}+\frac{\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}e^{-aN_{1}}\left(\frac{2}{N_{1}}+a\right),$ (14)
@@
$M_{\mathrm{PBH}}\sim 3\times 10^{48}\mathrm{g}\times e^{-2N_{1}}.$ (15)
Therefore, to have PBH as all of the dark matter in the present universe, i.e., $f_{\mathrm{PBH}}=1$ is not yet constraint from observation, $N_{1}=30\sim 40$ is needed  [^4] See reviews on the observational constraints on PBH as dark matter [61, 63, 92].
When the ratio of the power spectrum value at the peak scale and the CMB scale to be ${\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}/{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}\sim 10^{7}$ is fixed to generate a sizable abundance of primordial black holes today, i.e. $f_{\mathrm{PBH}}\sim 1$ [61, 62, 63], if the coupling constant $\xi$ is larger than 3/163/163 / 16 ($a=3$), the modifications from the blue-tilted component to the $n_{s}$ and $\alpha_{s}$ at CMB scale is exponentially suppressed, effectively decoupling small-scale features from large-scale observables. In this limit, the large-scale power spectrum reduces to that of the single-field $R^{2}$ inflation, with $\alpha_{s}\lesssim-0.0013$, $n_{s}\lesssim 0.95$ for $N_{1}\lesssim 40$. When $e^{-aN_{1}}$ is of order $10^{-7}$ or less, the modification to the large-scale spectrum comes from the small-scale enhancement can be significant.
@@
$M/M_{\rm pl}$ $1.35\times 10^{-5}$ $\ 1.78\times 10^{-5}$ $3.5\times 10^{-5}$
$m/M$ 1/61/61 / 6 1/6\ 1/6\ 1 / 6 1/61/61 / 6
$\xi$ 0.05300.05300.0530 0.06890.06890.0689 0.11630.11630.1163
$\chi_{0}/M$ 0.280.280.28 0.390.390.39 0.390.390.39
@@
$n_{s}(0.05\mathrm{Mpc}^{-1})$ 0.9750.9750.975 0.9730.9730.973 0.9730.9730.973
$n_{s}(0.002\mathrm{Mpc}^{-1})$ 0.9660.9660.966 0.9550.9550.955 0.9120.9120.912
$\alpha_{s}(0.05\mathrm{Mpc}^{-1})$ 0.0080.0080.008 0.0140.0140.014 0.0550.0550.055555Excluded at more than $2\sigma$ level.
@@
$f_{\mathrm{PBH}}$ 1$\sim$ 1∼ 1 1$\sim$ 1∼ 1 1$\ll$ 1≪ 1
Table 1: The benchmark parameters and the observables. The primordial power spectrum is normalized by the Planck 2018 result $\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=\mathcal{P}_{\mathcal{R}}(0.05\mathrm{Mpc}^{-1})=2.1\times 10^{-9}$. A joint P-ACT-LB analysis reported a spectral index $n_{s}(0.05{\rm Mpc}^{-1})=0.9743\pm 0.0034$. The total number of efolds $N$ of the inflation in our model can be tuned from 506050$\sim$ 6050 ∼ 60 without affecting the IR by adjusting $\lambda$, thus not listed in the table.
[…]
IV Numerical results

### v3

# Primordial black holes save $R^{2}$ Inflation

@@ 1 Introduction
However, a newly released result on the primordial power spectrum from the Atacama Cosmology Telescope (ACT) DR6 [79, 17] reported a slightly larger spectral index. Combined with Planck data at large-to-intermediate scales (denoted as P-ACT) and together with CMB lensing and BAO data from DESI [3] (denoted as P-ACT-LB), the analysis yields a new constraint on the spectral index: $n_{s}=0.9743\pm 0.0034$ (68% CL) at the CMB pivot scale $k_{\mathrm{CMB}}=0.05~\mathrm{Mpc}^{-1}$ [79]. This result appears to disfavor the $R^{2}$ inflation as illustrated in figure 5. In addition, the error bar for the running of the spectral index $\alpha_{s}$ is tightened in the P-ACT-LB result, which suggests a slight positive running of the spectral index $\alpha_{s}=0.0062\pm 0.0052$ (68% CL) [17]. The constraint on $\alpha_{s}$ is not consistent with most of the single-field models which involve a convex potential, including the $R^{2}$ model. Several works [12, 76, 25, 71, 45, 49, 114, 4, 39, 41] discussed extensions to the single-field $R^{2}$ model that may resolve this inconsistency.
@@
A crucial point in our model is the presence of a new boson field $\chi$ which has an $Z_{2}$ invariant potential where the $\chi$ is odd under the discrete symmetry. We identify it with the $Z_{4}$ symmetry whose breaking generates the Majorana masses $M_{i}$ for the right-handed neutrinos $N_{Ri}~(i=1-3)$  [69]. Here, the $\chi$ and $N_{Ri}$ have the $Z_{4}$ charges +2+2 and 1-1, respectively. This $Z_{4}$ is free from the Dai-Freed anomaly [29] and hence it is a perfect symmetry [109, 115] in the standard model. The breaking scale of the $Z_{4}$ is $\langle\chi\rangle\simeq 10^{12}$ GeV which gives the Majorana masses $M_{i}\sim g_{i}\langle\chi\rangle$. Those masses are very much consistent with the observed small masses for the neutrinos [86, 112, 113, 44].
The remainder of this paper is organized as follows. In section 2, we describe the setup and analyze the field dynamics. Section 3 presents the semi-analytical treatment of the enhancement in the curvature perturbations, and its implications for CMB observables such as the spectral index $n_{s}$ , its running $\alpha_{s}$ , and the tensor-to-scalar ratio $r$, as well as for the formation of primordial black holes (PBHs). section 4 compares our numerical results with the latest CMB measurements, highlighting the parameter regions consistent with current constraints. Finally, section 5 provides a summary and discussion of our main conclusions.
2 The model
We consider the action in which $R^{2}$ gravity is non-minimally coupled to a scalar field $\chi$ [94, 54, 25, 24, 108],
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg[\frac{M_{\mathrm{pl}}^{2}}{2}f(R,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)\bigg],$ (2.1)
@@ 2 The model
where $v=\pm m/\sqrt{\lambda}$ is the vacuum expectation value of $\chi$, and $m$ is the mass of $\chi$. Notice that the above action is invariant under the $Z_{4}$ symmetry under which the $\chi$ transforms to $\chi$. The presence of $|\chi|$ is not a serious problem, since we consider it to be the low-energy action in an effective field theory (EFT). In fact, the $|\chi|$ dependent terms often appear at the EFT level when such a boson field couples to non-abelian gauge theories. (See [50] for example.)
It is well-known that the $R^{2}$ term introduces a scalar degree of freedom to the theory. Introducing an auxiliary field $\psi$, the action (2.1) is equivalent to,
$S_{\psi}=\int d^{4}x\sqrt{-g}\ \bigg[\frac{M_{\mathrm{pl}}^{2}}{2}RF(\psi,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)-U(\psi)\bigg],$ (2.2)
@@
$F(\psi,\chi)$ $=1+\frac{\psi}{6M^{2}}-\frac{\xi}{M_{\mathrm{pl}}^{2}}(|\chi|-\chi_{0})^{2}$ (2.3)
$U(\psi)$ $=\frac{M_{\mathrm{pl}}^{2}\psi^{2}}{12M^{2}}$ (2.4)
To bring the above action into the Einstein frame, we perform the conformal transformation $\tilde{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ [30] for the choice of $\Omega^{2}=F$, where the $\Omega^{2}$ is the conformal factor, and a tilde represents quantities in the Einstein frame. Under transformation, the action (2.1) can be written in the Einstein frame,
$S_{\mathrm{E}}=\int d^{4}x\sqrt{-\tilde{g}}\left[\frac{M_{\mathrm{pl}}^{2}}{2}\tilde{R}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}F^{-1}\tilde{g}^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-U(\phi,\chi)\right],$ (2.5)
@@
    Given the large initial value of $\phi$, the spectator field $\chi$ acquires an effective mass approximately given by $m_{\chi}^{2}\approx 3M^{2}\xi(1+\mathcal{O}(F(\phi)^{-1}))$. We assume that the initial value of $\chi$ is close to its potential minimum $\pm\chi_{0}$ at the beginning of inflation. During this stage, when $m_{\chi}^{2}>9H^{2}/4\approx 9M^{2}/16$, or equivalently $\xi>3/16$, $\chi$ is considered a heavy field and behaves as an underdamped oscillator. In contrast, when $m_{\chi}^{2}<9H^{2}/4$, $\chi$ is light and behaves as an overdamped oscillator; note that in this case, $\phi$ remains the lightest field. In each of the cases, the scalar field $\chi$ is settled down to one of its potential minimum ($\chi_{0}$ or $\chi_{0}$, $\chi_{0}$ is taken for the computations in the following section) before this stage of inflation reaches an end.
@@
    $U(\phi,\chi)\approx(\lambda/4)(\chi_{0}^{2}-v^{2})^{2}\approx\frac{m^{4}}{4\lambda}-\frac{1}{2}m^{2}\chi^{2}$ (2.7)
    As shown above, the boson field $\chi$ becomes unstable $(m_{\chi}^{2}\approx-m^{2}<0)$. Therefore, the inflaton dynamics begin to turn toward the $\pm\chi$ direction. It is worth mentioning that $\chi$ field starting from a initial condition away from 0 steps down the potential during the first stage of the $R^{2}$ inflation and hence the $Z_{4}$ symmetry is already broken for the background trajectory, or more precisely,inflation selects a vacuum branch during the early stage (see in figure 1). When fixing $|\chi_{0}|^{2}\gg\langle\delta\chi^{2}\rangle$  [95, 13], the overproduction of domain walls (DWs) can be avoided, which usually happens in hybrid inflation models with a discrete $Z_{2}$ symmetry.
  • St-2: Effective single field $\chi$ inflation. The second stage of inflation starts along $\chi$ direction from $\pm\chi_{0}$ (the sign will depend on the initial condition for $\chi$), with the heavy field $\phi$ stabilized in its potential minimum. We may observe that $\phi$ is still evolving during this stage, but it is simply due to the varying $\chi$ causing the movement of the potential minimum. The time evolution of $\phi$ is not violating the effective single field slow condition, since the isocurvature is heavy and decays out very rapidly. The second stage of inflation ends as an effective single field inflation of $\chi$. To have a sizable number of efolds during St-2, $v\gg\chi_{0}$ is required. At the same time, eternal inflation may happen during waterfall region due to the flatness of the potential along the background trajectory, which gives another constraint on $\chi_{0}$. To avoid eternal inflation, the classical displacement of the field within one Hubble time should be much larger than the quantum fluctuation, requiring
    $\Delta\chi_{\mathrm{cl}}\equiv\frac{\left|\dot{\chi}\right|}{H}\gg\sqrt{\langle\delta\chi^{2}\rangle},$ (2.8)
    where the typical size of inflaton quantum $\sqrt{\langle\delta\chi^{2}\rangle}\sim H/(2\pi)$ in the superhorizon limit.
    Combining with conditions to avoid DW overproduction and eternal inflation, the constraints for the parameters is given
    $\frac{m}{\sqrt{\lambda}}\gg|\chi_{0}|\gg\max\left\{\frac{M}{4\pi},\frac{m^{4}}{16\sqrt{3}\pi\lambda^{3/2}}\right\}$ (2.9)
@@
Thus, by varying the coupling constant $\xi$ from 0 to values $\geq 3/16$, the model allows the contribution from $\chi$ to the adiabatic power spectrum to evolve from nearly scale-invariant ($k^{0}$) to steeply blue-tilted ($k^{3}$), as explained detailedly in the following section (also shown in [108]).
@@
3 Semi-analytical power spectrum
@@
For simplicity, we assume the transition between the two inflationary phases occurs instantaneously. A schematic illustration of the background and perturbed trajectories is shown in figure 2.
@@
$\mathcal{R}=\delta N=\delta N_{1}+\delta N_{2},$ (3.1)
where $N_{1}\equiv n_{*}-n_{k}$ and $N_{2}\equiv n_{\mathrm{f}}-n_{*}$ are the total number of e-folds during the first and second stages of inflation, respectively. Here we introduce a new time variable $(t)=\int_{t_{i}}^{t}Hdt, nn_{*}$ and $n_{\mathrm{f}}$ represent the number of e-folds respect to an arbitrarily set initial time $t_{i}$ at the transition and at the end of inflation.
@@
If we conduct the $\delta N$ calculation nonlinearly, one may observe a large deviation of the probability distribution function of $\mathcal{R}$ from Gaussianity, as observed already in single field inflation models [16, 15, 93, 67, 106]. Such non-Gaussian features are potentially detectable by upcoming observations. We leave this for future work.
@@
$\delta N=\left(\frac{\partial N_{1}}{\partial\phi}\delta\phi_{k}\right)_{n_{k}}+\left(\frac{\partial N_{2}}{\partial\chi}\delta\chi_{k}\right)_{n_{\star}},$ (3.2)
@@
$\delta\chi_{k}(n_{\star})\approx\delta\chi_{k}(n_{k})e^{-a(n_{\star}-n_{k})/2},$ (3.3)
@@
$a$ $\equiv\mathrm{Re}\left(3-3\sqrt{1-\frac{16m^{2}_{\chi}}{9M^{2}}}\right)$ (3.4)
@@
during the St-1, which equals the maximum $3$ for the heavy $\chi$ case and smaller when decreasing $\xi$.
Figure 3: The $n_{s}-\alpha_{s}$ relation shown in eq. (3.6) for different $M_{\mathrm{PBH}}$s together with the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [17]). To plot the figure, we use $\mathcal{P}_{\mathcal{R}}^{\mathrm{max}}/\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=10^{7}$ in order to have $f_{\mathrm{PBH}}=1$. The hollow black dots refer to cases that $a=3$, or equivalently $\xi\geq 3/16$ when the small-scale amplification is well isolated from the large scales.
@@
$\mathcal{P}_{\mathcal{R}}(k)=\mathcal{P}_{\mathcal{R}}^{R^{2}}(k)+\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}\left(\frac{k}{k_{1}}\right)^{a}.$ (3.5)
@@
$n_{s}\approx 1-\frac{2}{N_{1}}+\frac{\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}e^{-aN_{1}}\left(\frac{2}{N_{1}}+a\right),$ (3.6)
@@
$M_{\mathrm{PBH}}\sim 10^{18}\mathrm{g}\times e^{-2(N_{1}-35)}.$ (3.7)
Therefore, to have PBH as all of the dark matter in the present universe, i.e., $f_{\mathrm{PBH}}=1$ is not yet constrained by observations, $N_{1}=30\sim 40$ is needed  [^4] See reviews on the observational constraints on PBH as dark matter [21, 100, 20].
When the ratio of the power spectrum value at the peak scale and the CMB scale to be ${\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}/{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}\sim 10^{7}$ is fixed to generate a sizable abundance of primordial black holes today, i.e. $f_{\mathrm{PBH}}\sim 1$ [21, 47, 100], if the coupling constant $\xi$ is larger than $3/16$ ($a=3$), the modifications from the blue-tilted component to the $n_{s}$ and $\alpha_{s}$ at CMB scale is exponentially suppressed, effectively decoupling small-scale features from large-scale observables. In this limit, the large-scale power spectrum reduces to that of the single-field $R^{2}$ inflation, with $\alpha_{s}\lesssim-0.0013$, $n_{s}\lesssim 0.95$ for $N_{1}\lesssim 40$. When $e^{-aN_{1}}$ is of order $10^{-7}$ or less, the modification to the large-scale spectrum comes from the small-scale enhancement can be significant.
@@
$M/M_{\rm pl}$ $1.35\times 10^{-5}$ $1.78\times 10^{-5}\$$3.5\times 10^{-5}$
$m/M$ $1/6$ 1/6\ 1/6\ $1/6$
$\xi$ $0.0530$ $0.0689$ $0.1163$
$\chi_{0}/M$ $0.28$ $0.39$ $0.39$
$\lambda$ $4.7\times 10^{-12}$ $2.9\times 10^{-12}$ $6.8\times 10^{-12}$
@@
$n_{s}(0.05\mathrm{Mpc}^{-1})$ $0.975$ $0.973$ $0.973$
$n_{s}(0.002\mathrm{Mpc}^{-1})$ $0.966$ $0.955$ $0.912$
$\alpha_{s}(0.05\mathrm{Mpc}^{-1})$ $0.008$ $0.014$ $0.055,^{(a)}$
@@
$f_{\mathrm{PBH}}$ $\sim 1$ $\sim 1$ $\ll 1$
Table 1: The benchmark parameters and the observables. The primordial power spectrum is normalized by the Planck 2018 result $\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=\mathcal{P}_{\mathcal{R}}(0.05\mathrm{Mpc}^{-1})=2.1\times 10^{-9}$. A joint P-ACT-LB analysis reported a spectral index $n_{s}(0.05{\rm Mpc}^{-1})=0.9743\pm 0.0034$. The total number of efolds NN of the inflation in our model can be tuned from $50\sim 60$ without affecting the IR by adjusting $\lambda$, thus not listed in the table. Notes: Excluded at more than $2\sigma$ level.
Figure 4: The primordial power spectrum evaluated at the end of inflation shown alongside the (future) observational constraints. The black dashed line shows the result for $\xi>3/16$, whose power spectrum converges to the effective single-field $R^{2}$ inflation. The curves are obtained by numerically solving for the primordial power spectrum by the Transport method using the model parameters from Table 1, Case 2 except for a varying $\xi$. From the darkest to the lightest solid curves, we take $\xi=0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (Case 2), $0.0685$. The constraint plot comes from [66], including CMB constraints (red) from Planck [8], Lyman-$\alpha$ forest constraints [116] (blue), $\mu$-distortion constraints (green) from FIRAS [59], and PIXIE (future) [2, 74], and gravitational wave constraints (purple) from PTAs [7, 11, 111], SKA [65] (future), LISA [10] (future) and BBO [51] (future).
4 Numerical results